AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 299     Entry time: Sun Feb 27 08:00:23 2022
Author: RSS TD 
Subject: Current status 

Ring current ~7.5e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.44 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 9-xy-2022-02-27_09-00-57.png  274 kB  | Hide | Hide all
9-xy-2022-02-27_09-00-57.png
Attachment 2: 9-ey-2022-02-27_09-00-03.png  167 kB  | Hide | Hide all
9-ey-2022-02-27_09-00-03.png
Attachment 3: 9-ex-2022-02-27_08-59-48.png  167 kB  | Hide | Hide all
9-ex-2022-02-27_08-59-48.png
ELOG V3.1.4-unknown